# Revisiting Non-Acyclic GFlowNets in Discrete Environments

### v1

@@ Appendix A Proofs
Thus, the optimal policy $\pi^{\star}$ that maximizes $V^{\pi}_{\lambda=1}(s_{0})$ is the one that minimizes $\operatorname{KL}(\mathsf{P}_{\mathcal{T}}^{\pi}|\mathcal{P})$. By Proposition 3.8, there exists a unique forward policy $\mathcal{P}_{\mathrm{F}}$ that induces the same trajectory distribution as $\mathcal{P}_{\mathrm{B}}$, which is equivalent to achieving zero KL-divergence. Thus, $\pi^{\star}$ coincides with $\mathcal{P}_{\mathrm{F}}$, and we conclude the statement by the uniqueness of the solution (see Lemma A.1). To apply Lemma A.1, without loss of generality, we can assume that the GFlowNet reward function $\mathcal{R}$ is normalized, i.e., $\mathcal{Z}=1$ and $\log\mathcal{R}(x)\leq 0$. Indeed, since a terminating transition $x\to s_{f}$ is always visited exactly once, it is equivalent to subtracting $\log\mathcal{Z}$ from all terminal rewards, which does not change the optimal policy and modifies all values by the same constant.
@@
$\begin{split}\mathcal{F}(s\to s^{\prime})=\log\mathcal{F}(s^{\prime})+\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})&=\log\sum_{s^{\prime\prime}\in\operatorname{out}(s^{\prime})}\mathcal{F}(s^{\prime}\to s^{\prime\prime})+\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})\\ &=\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})+\log\sum_{s^{\prime\prime}\in\operatorname{out}(s^{\prime})}\exp\mathopen{}\mathclose{{}\left(\log\mathcal{F}(s^{\prime}\to s^{\prime\prime})}\right).\end{split}$
@@ B.1 Training Policy and Flow Weighting
This implies that on-policy training tries to minimize the sum of squared state flows rather than the sum of state flows. This happens due to the fact that the trajectory distribution that is used to collect data for training (induced by $\mathcal{P}_{\mathrm{F}}$ in this case) visits certain states more often then others, thus a weight is given to the flow in each state equal to the expected number of visits. However, if $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ is fixed to be uniform over $\mathcal{S}\setminus\{s_{0},s_{f}\}$ (see Section 4 and Appendix B.3), this issue can be circumvented by applying flow regularizer only in the first state of each sampled trajectory. Then, equal weight will be given to $\mathcal{F}_{\theta}(s)$ in each state in the expected loss, thus we will be minimizing the sum of state flows. However, in our experiments we noticed that this does not significantly influence the results, thus we leave exploring this phenomenon as a further research direction.
@@ B.2 Loss Scaling and Stability
In this section, we provide a more detailed explanation of our scaling hypothesis (see Section 4). Let us consider a GFlowNet that learns $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$. Since these quantities are predicted by a neural network, a standard way is to make it predict logits for the forward policy, logits for the backward policy, and logarithm of the state flow. Flow functions are always positive, thus predicting them in log scale is a natural approach (Bengio et al., 2021, 2023). Then, for any transition $s\to s^{\prime}$, define two quantities:
@@
The first is difference between predicted logarithms of the flows in the forward and backward direction $\log\mathcal{F}_{F}-\log\mathcal{F}_{B}$, while the second is difference between predicted flows in the forward and backward direction $\mathcal{F}_{F}-\mathcal{F}_{B}$. Then, the standard DB loss (4) is
@@
However, for both losses one can either replace $\Delta_{\log\mathcal{F}}$ with $\Delta_{\mathcal{F}}$ or the other way around. For visualization, let us fix the predicted log backward flow $\mathcal{F}_{B}$ to be, e.g., $1$, and plot the losses with respect to the varying value of the predicted log forward flow $\mathcal{F}_{F}$. The plots are presented in Figure 3. One can note that as argument $\log\mathcal{F}_{F}$ decreases, both losses in $\Delta_{\mathcal{F}}$ scale quickly plato, thus their derivative goes to zero. From the optimization perspective this means that when the predicted log flow needs to be increased, the gradient step will be very small since the derivative of the loss is almost zero. On the other hand, when the predicted log flow needs to be decreased, the gradient step will be larger since losses have much higher derivatives in the corresponding regions. In combination with Proposition 3.12, this gives a possible explanation to stability of $\Delta_{\mathcal{F}}$ scale losses: they are biased towards underestimation of the flows, and, as a result, biased towards solutions with smaller expected trajectory length. We note that the same reasoning can be applied to stable flow matching loss proposed in (Brunswic et al., 2024) since it also operates with differences between flows in $\Delta_{\mathcal{F}}$ scale.
@@ B.3 Fixed $\mathcal{P}_{\mathrm{B}}$ and Trainable $\mathcal{P}_{\mathrm{B}}$
In non-acyclic environments, $s_{0}$ and $s_{f}$ generally are fictive states that do not correspond to any object. Then $\mathcal{P}_{\mathrm{F}}(s_{f}\mid s)$ corresponds to probability to terminate a trajectory in state $s$, while $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ corresponds to the probability that a trajectory starts in the state $s$. Thus, the choice of $\operatorname{out}(s_{0})$ is crucial in the design of the environment. If this set is large, e.g., coincides with $\in\mathcal{S}\setminus\{s_{0},s_{f}\}$, one has to fix $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ to some distribution, e.g. uniform, otherwise learning becomes intractable. However, in this case $\mathcal{P}_{\mathrm{B}}(s_{0}\mid s)$ has to be trainable, otherwise it may be impossible to satisfy detailed balance conditions for transitions $s_{0}\to s$.
@@ B.4 Solving Small Environments Exactly
Interestingly, the system (17) can also be explained from Markov Chain perspective. Let us take the graph $\mathcal{G}$ with reversed edges, add a loop from $s_{0}$ to itself, and use $\mathcal{P}_{\mathrm{B}}$ to define a Markov Chain: $P(s_{0}\mid s_{0})=1$, $P(s\mid s^{\prime})=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ if there is an edge $s\to s^{\prime}$, and $P(s\mid s^{\prime})=0$ otherwise. It will be an absorbing Markov Chain, with an only absorbing state $s_{0}$ since it is reachable from any other state by Assumption 3.1. Its transition matrix can be written in the following way:
@@
where $I-Q$ is always invertible (Kemeny et al., 1969, Theorem 3.2.1). One can note that normalized flows $\mathcal{F}(s)/\mathcal{F}(s_{f})$ coincide with the expected number of visits to $s$ when starting from $s_{f}$, thus coincide with the row of matrix $N$ corresponding to $s_{f}$. Finally, notice that $(I-Q)$ coincides with the transposed matrix of the truncated system (17) (with the exception of the variable and the equation corresponding to $s_{0}$), thus such system has a unique solution $\mathcal{F}(s_{f})(I-Q)^{-T}e_{s_{f}}=\mathcal{F}(s_{f})N^{T}e_{s_{f}}$, where $e_{s_{f}}$ is a vector of size $|\mathcal{S}|-1$ that has $1$ on the position corresponding to $s_{f}$ and 00 on all others. Variable corresponding to $s_{0}$ should be handled separately, but it is easy to see ^$\hat{\mathcal{F}}(s_{0})=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}){\mathcal{F}}(s^{\prime})=\mathcal{F}(s_{0})$.
@@ C.1 Loss Choice
While (Brunswic et al., 2024) used the original flow matching loss (Bengio et al., 2021) for experimental evaluation, it was previously shown to be less computationally efficient and provide slower convergence than other GFlowNet losses (Malkin et al., 2023a; Madan et al., 2023) in acyclic case, so we carry out experimental evaluation with the more broadly employed detailed balance loss (Bengio et al., 2023). Moreover, flow matching loss does not admit explicit parameterization of a backward policy, as well as training with fixed backward policies, thus not allowing us to study some of the phenomena we explore in the experiments.
@@ C.2 Hypergrids
where $\pi(x)$ is empirical distribution of last $2\cdot 10^{5}$ samples seen in training (endpoints of trajectories sampled from $\mathcal{P}_{\mathrm{F}}$).
All models are parameterized by MLP with 2 hidden layers and 256 hidden size, which accept a one-hot encoding of $s$ as input. $\mathcal{F}_{\theta}(s),\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s,\theta),\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime},\theta)$ share the same backbone, with different linear heads predicting logarithm of the state flow, logits of the forward policy and logits of the backward policy. In case of the fixed $\mathcal{P}_{\mathrm{B}}$, $s_{\text{init}}$ corresponds to the center of the grid and we take $\varepsilon=10^{-8}$ (see Appendix B.3).
We train all models on-policy. We use Adam optimizer with a learning rate of $10^{-3}$ and a batch size 16 (number of trajectories sampled at each training step). For $\log\mathcal{Z}_{\theta}$ we use a larger learning rate of $10^{-2}$ (see (Malkin et al., 2022)). All models are trained until $2\cdot 10^{6}$ trajectories are sampled, and the empirical sample distribution $\pi(x)$ is computed over the last $2\cdot 10^{5}$ samples seen in training. For SDB we set $\varepsilon=1.0$ and $\eta=10^{-3}$. We found that using larger values of $\eta$ can lead to smaller expected trajectory length, but also significantly interfere with sampling fidelity of the learned GFlowNet, thus we opt for these values in our experiments.
@@ C.3 Permutations
All models are parameterized by MLP with 2 hidden layers and 128 hidden size, which accept a one-hot encoding of $s$ as input. $\mathcal{F}_{\theta}(s),\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s,\theta),\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime},\theta)$ share the same backbone, with different linear heads predicting logarithm of the state flow, logits of the forward policy and logits of the backward policy. In case of the fixed $\mathcal{P}_{\mathrm{B}}$, $s_{\text{init}}$ corresponds to the permutation $(n,n-1,\dots,2,1)$ and we take $\varepsilon=10^{-8}$ (see Appendix B.3).
We train all models on-policy. We use Adam optimizer with a learning rate of $10^{-3}$ and a batch size 512 (number of trajectories sampled at each training step). We found that using small batch sizes can significantly hinder training stability for this environment; thus, we opt for a larger value. All models are trained for $10^{5}$ iterations. For $\log\mathcal{Z}_{\theta}$ we use a larger learning rate of $10^{-2}$ (see (Malkin et al., 2022)). Empirical distribution $\hat{C}(k)$ is computed over the last $10^{5}$ samples seen in training. For SDB we set $\varepsilon=1.0$ and $\eta=10^{-3}$. We found that using larger values of $\eta$ can potentially lead to smaller expected trajectory length, but also significantly interfere with sampling fidelity of the learned GFlowNet, thus we opt for these values in our experiments.
@@
where $\mathbb{E}[\mathcal{R}(x)]=\sum_{x\in\mathcal{X}}\mathcal{R}(x)\frac{\mathcal{R}(x)}{\mathcal{Z}}$.
@@ C.3.1 Reward Distribution Properties
We define the GFlowNet reward as $\mathcal{R}(s)=\exp(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\})$. We are interested in true values of three quantities:
@@
since choosing a permutation with $k$ fixed points coincides with choosing $k$ positions for fixed points, and permuting the remaining elements such that there are no fixed points among them. So let us start with derivation of $D(0,n)$. Denote $S_{i}$ to be the set of permutations on $n$ elements that has a fixed point on position $i$. Then, by inclusion-exclusion principle we have

### v2

@@ Appendix A Proofs
Thus, an optimal policy $\pi^{\star}$ that maximizes $V^{\pi}_{\lambda=1}(s_{0})$ is the one that minimizes $\operatorname{KL}(\mathsf{P}_{\mathcal{T}}^{\pi}|\mathcal{P})$. By Lemma 3.4, the expected trajectory length of any optimal policy $\pi^{\star}$ that matches $\mathcal{P}$ is finite. Thus, by Proposition 3.8, there exists a unique forward policy $\mathcal{P}_{\mathrm{F}}$ that induces the same trajectory distribution as $\mathcal{P}_{\mathrm{B}}$, which is equivalent to achieving zero KL-divergence. Thus, $\pi^{\star}$ coincides with $\mathcal{P}_{\mathrm{F}}$, and we conclude the statement by the uniqueness of the solution (see Lemma A.1). To apply Lemma A.1, without loss of generality, we can assume that the GFlowNet reward function $\mathcal{R}$ is normalized, i.e., $\mathcal{Z}=1$ and $\log\mathcal{R}(x)<0$. Indeed, since a terminating transition $x\to s_{f}$ is always visited exactly once, it is equivalent to subtracting $\log\mathcal{Z}$ from all terminal rewards, which does not change the optimal policy and modifies all values by the same constant.
@@
$\begin{split}\log\mathcal{F}(s\to s^{\prime})=\log\mathcal{F}(s^{\prime})+\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})&=\log\sum_{s^{\prime\prime}\in\operatorname{out}(s^{\prime})}\mathcal{F}(s^{\prime}\to s^{\prime\prime})+\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})\\ &=\log\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})+\log\sum_{s^{\prime\prime}\in\operatorname{out}(s^{\prime})}\exp\mathopen{}\mathclose{{}\left(\log\mathcal{F}(s^{\prime}\to s^{\prime\prime})}\right).\end{split}$
@@ B.1 Training Policy and Flow Weighting
This implies that on-policy training tries to minimize the sum of squared state flows rather than the sum of state flows. This happens due to the fact that the trajectory distribution that is used to collect data for training (induced by $\mathcal{P}_{\mathrm{F}}$ in this case) visits certain states more often than others, thus a weight is given to the flow in each state equal to the expected number of visits. However, if $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ is fixed to be uniform over $\mathcal{S}\setminus\{s_{0},s_{f}\}$ (see Section 4 and Appendix B.3), this issue can be circumvented by applying flow regularizer only in the first state of each sampled trajectory. Then, equal weight will be given to $\mathcal{F}_{\theta}(s)$ in each state in the expected loss, thus we will be minimizing the sum of state flows. However, in our experiments we noticed that this does not significantly influence the results, so we leave exploring this phenomenon as further research direction.
@@ B.2 Loss Scaling and Stability
In this section, we provide a more detailed explanation of our scaling hypothesis (see Section 4). Let us consider a GFlowNet that learns $\mathcal{F}$, $\mathcal{P}_{\mathrm{F}}$ and $\mathcal{P}_{\mathrm{B}}$. Since these quantities are predicted by a neural network, a standard way is to make it predict logits for the forward policy, logits for the backward policy, and the logarithm of the state flow. Flow functions are always positive, thus predicting them in log scale is a natural approach (Bengio et al., 2021, 2023). Then, for any transition $s\to s^{\prime}$, define two quantities:
@@
The first is the difference between the predicted logarithms of the flows in the forward and backward direction $\log\mathcal{F}_{F}-\log\mathcal{F}_{B}$, while the second is the difference between predicted flows in the forward and backward direction $\mathcal{F}_{F}-\mathcal{F}_{B}$. Then, the standard DB loss (4) is
@@
However, for both losses, one can either replace $\Delta_{\log\mathcal{F}}$ with $\Delta_{\mathcal{F}}$ or the other way around. For visualization, let us fix the predicted log backward flow $\mathcal{F}_{B}$ to be, e.g., $1$, and plot the losses with respect to the varying value of the predicted log forward flow $\mathcal{F}_{F}$. The plots are presented in Figure 3. One can note that as argument $\log\mathcal{F}_{F}$ decreases, both losses in $\Delta_{\mathcal{F}}$ scale quickly plateau, thus their derivative goes to zero. From the optimization perspective, this means that when the predicted log flow needs to be increased, the gradient step will be very small since the derivative of the loss is almost zero. On the other hand, when the predicted log flow needs to be decreased, the gradient step will be larger since losses have much higher derivatives in the corresponding regions. In combination with Proposition 3.12, this gives a possible explanation to the stability of $\Delta_{\mathcal{F}}$ scale losses: they are biased towards underestimation of the flows, and, as a result, biased towards solutions with smaller expected trajectory length. We note that the same reasoning can be applied to the stable flow matching loss proposed in (Brunswic et al., 2024) since it also operates with differences between flows in $\Delta_{\mathcal{F}}$ scale.
@@ B.3 Fixed $\mathcal{P}_{\mathrm{B}}$ and Trainable $\mathcal{P}_{\mathrm{B}}$
In non-acyclic environments, $s_{0}$ and $s_{f}$ generally are fictive states that do not correspond to any object. Then $\mathcal{P}_{\mathrm{F}}(s_{f}\mid s)$ corresponds to the probability to terminate a trajectory in state $s$, while $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ corresponds to the probability that a trajectory starts in the state $s$. Thus, the choice of $\operatorname{out}(s_{0})$ is crucial in the design of the environment. If this set is large, e.g., coincides with $\in\mathcal{S}\setminus\{s_{0},s_{f}\}$, one has to fix $\mathcal{P}_{\mathrm{F}}(s\mid s_{0})$ to some distribution, e.g. uniform, otherwise learning becomes intractable. However, in this case $\mathcal{P}_{\mathrm{B}}(s_{0}\mid s)$ has to be trainable, otherwise, it may be impossible to satisfy the detailed balance conditions for transitions $s_{0}\to s$.
@@ B.4 Solving Small Environments Exactly
Interestingly, the system (17) can also be explained from Markov Chain perspective. Let us take the graph $\mathcal{G}$ with reversed edges, add a loop from $s_{0}$ to itself, and use $\mathcal{P}_{\mathrm{B}}$ to define a Markov Chain with the following transition matrix: $P(s_{0}\mid s_{0})=1$, $P(s\mid s^{\prime})=\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime})$ if there is an edge $s\to s^{\prime}$, and $P(s\mid s^{\prime})=0$ otherwise. It will be an absorbing Markov Chain, with an only absorbing state $s_{0}$ since it is reachable from any other state by Assumption 3.1. The transition matrix can be written in the following way:
@@
where $I-Q$ is always invertible (Kemeny & Snell, 1969, Theorem 3.2.1). One can note that normalized flows $\mathcal{F}(s)/\mathcal{F}(s_{f})$ coincide with the expected number of visits to $s$ when starting from $s_{f}$, thus coincide with the row of matrix $N$ corresponding to $s_{f}$. Finally, notice that $(I-Q)$ coincides with the transposed matrix of the truncated system (17) (with the exception of the variable and the equation corresponding to $s_{0}$), thus such system has a unique solution $\mathcal{F}(s_{f})(I-Q)^{-T}e_{s_{f}}=\mathcal{F}(s_{f})N^{T}e_{s_{f}}$, where $e_{s_{f}}$ is a vector of size $|\mathcal{S}|-1$ that has $1$ on the position corresponding to $s_{f}$ and 00 on all others. The variable corresponding to $s_{0}$ should be handled separately, but it is easy to see ^$\hat{\mathcal{F}}(s_{0})=\sum_{s^{\prime}\in\operatorname{out}(s)}\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime}){\mathcal{F}}(s^{\prime})=\mathcal{F}(s_{0})$.
@@ C.1 Loss Choice
While (Brunswic et al., 2024) used the original flow matching loss (Bengio et al., 2021) for experimental evaluation, it was previously shown to be less computationally efficient and provide slower convergence than other GFlowNet losses (Malkin et al., 2023; Madan et al., 2023) in the acyclic case, so we carry out experimental evaluation with the more broadly employed detailed balance loss (Bengio et al., 2023). Moreover, flow matching loss does not admit explicit parameterization of a backward policy, as well as training with fixed backward policies, thus not allowing us to study some of the phenomena we explore in the experiments.
@@ C.2 Hypergrids
where $\pi(x)$ is the empirical distribution of last $2\cdot 10^{5}$ samples seen in training (endpoints of trajectories sampled from $\mathcal{P}_{\mathrm{F}}$).
All models are parameterized by MLP with 2 hidden layers and 256 hidden size, which accept a one-hot encoding of $s$ as input. $\mathcal{F}_{\theta}(s),\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s,\theta),\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime},\theta)$ share the same backbone, with different linear heads predicting the logarithm of the state flow, the logits of the forward policy and the logits of the backward policy. In the case of the fixed $\mathcal{P}_{\mathrm{B}}$, $s_{\text{init}}$ corresponds to the center of the grid, and we take $\varepsilon=10^{-8}$ (see Appendix B.3).
We train all models on-policy. We use Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 16 (number of trajectories sampled at each training step). For $\log\mathcal{Z}_{\theta}$ we use a larger learning rate of $10^{-2}$ (see (Malkin et al., 2022)). All models are trained until $2\cdot 10^{6}$ trajectories are sampled, and the empirical sample distribution $\pi(x)$ is computed over the last $2\cdot 10^{5}$ samples seen in training. For SDB we set $\varepsilon=1.0$ and $\eta=10^{-3}$. We found that using larger values of $\eta$ can lead to smaller expected trajectory length, but also significantly interfere with the sampling fidelity of the learned GFlowNet, thus we opt for these values in our experiments.
@@ C.3 Permutations
All models are parameterized by MLP with 2 hidden layers and 128 hidden size, which accept a one-hot encoding of $s$ as input. $\mathcal{F}_{\theta}(s),\mathcal{P}_{\mathrm{F}}(s^{\prime}\mid s,\theta),\mathcal{P}_{\mathrm{B}}(s\mid s^{\prime},\theta)$ share the same backbone, with different linear heads predicting the logarithm of the state flow, the logits of the forward policy and the logits of the backward policy. In the case of the fixed $\mathcal{P}_{\mathrm{B}}$, $s_{\text{init}}$ corresponds to the permutation $(n,n-1,\dots,2,1)$, and we take $\varepsilon=10^{-8}$ (see Appendix B.3).
We train all models on-policy. We use Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 512 (number of trajectories sampled at each training step). We found that using small batch sizes can significantly hinder training stability for this environment; thus, we opt for a larger value. All models are trained for $10^{5}$ iterations. For $\log\mathcal{Z}_{\theta}$ we use a larger learning rate of $10^{-2}$ (see (Malkin et al., 2022)). To compute $\Delta\log\mathcal{Z}$ we take the average value of $\log\mathcal{Z}_{\theta}$ over the last 10 training checkpoints. Empirical distribution $\hat{C}(k)$ is computed over the last $10^{5}$ samples seen in training. For SDB we set $\varepsilon=1.0$ and $\eta=10^{-3}$. We found that using larger values of $\eta$ can lead to smaller expected trajectory length, but also significantly interfere with the sampling fidelity of the learned GFlowNet, thus we opt for these values in our experiments.
@@
where $\mathbb{E}[\mathcal{R}(x)]=\sum_{x\in\mathcal{X}}\mathcal{R}(x)\frac{\mathcal{R}(x)}{\mathcal{Z}}$. We compute mean reward over $10^{4}$ samples.
@@ C.3.1 Reward Distribution Properties
We define the GFlowNet reward as $\mathcal{R}(s)=\exp(\frac{1}{2}\sum_{k=1}^{n}\mathbb{I}\{s(k)=k\})$. We are interested in the true values of three quantities:
@@
since choosing a permutation with $k$ fixed points coincides with choosing $k$ positions for fixed points, and permuting the remaining elements such that there are no fixed points among them. So let us start with the derivation of $D(0,n)$. Denote $S_{i}$ to be the set of permutations on $n$ elements that has a fixed point on position $i$. Then, by the inclusion-exclusion principle, we have